# S2malloc: Statistically Secure Allocator for Use-After-Free Protection And More

### v1

@@ 4.4. Hardening heap canaries
4.5. Free block canaries (FBC)
@@
Initially in S2malloc, we create the memory pool using the mmap system call with the ANON flag and all allocated memories are set to zero in the Linux environment (Linux Foundation, 2023a). We take this advantage and use the zeros as the initial FBC with the following benefits:
@@ 4.6. Summary and comparison
Being an entropy-based allocator, S2malloc is inherently closer to this line of work (Novark and Berger, 2010; Silvestro et al., 2018; Liu et al., 2019) with a nearly identical set of heap exploitation protection features except UAF protection. S2malloc provides a much stronger security assurance in the presence of UAF vulnerabilities. In particular, S2malloc addresses the two entropy-loss cases (discussed in § 3.2 and § 3.3) with RIO (§ 4.2) and RBL (§ 4.3), respectively, and hence, providing much higher effectiveness on UAF mitigation. In addition, S2malloc is designed to actively monitor the integrity of the heap and watch for UAF attempts, including heap spraying practices that aim to prepare the heap data and layout for UAF exploits. S2malloc achieves this through a synergy of regular heap canaries (§ 4.4) and FBC (§ 4.5).
On the other hand, Table 1 also shows a sheer contrast between entropy-based allocators and UAF-mitigating allocators. Notably, although providing a theoretically complete mitigation guarantee toward UAF, UAF-mitigating allocators (Ainsworth and Jones, 2020; Wickman et al., 2021) significantly impair program efficiency and are hard to be deployed in time-sensitive use cases. In contrast, as will be presented in § 6, S2malloc incurs a significantly lower overhead that is typical for entropy-based allocators, making S2malloc practical and deployable on production systems if the end-user can tolerate a marginal chance of protection failure (less than 10% in the default setting of S2malloc, discussed in § 5).
5. Security Evaluation
@@
Figure 3. Parameterized security evaluation (x-axis: logarithmic block size/byte, y-axis: protection/attack success rate/%).
@@ 5.1. Formal analysis
For each of the scenarios above, we discuss the rates of whether the attacker adopts the heap spray technique or not separately. We introduce a total of four scenarios. We refer readers to Appendix D for a detailed explanation of them.
5.2. Parameterized protection rates
@@
2(a) shows how the protection rate changes with a different set of tunable parameters while 2(b) shows how the attack success rate changes using strategy S1. We observe that adopting a more secure setting increases the protection rate and decreases the attack success rates for both small and large blocks. However, as the block size increases, both rates decrease as it is less likely to overwrite either the target field or FBC. Both the attacker and the defender are likely not to succeed, and the rates are thus not summed up to one (in fact, not even close to one).
2(c) shows how the protection rate changes with the same set of tunable parameters while 2(d) shows how the attack success rate changes using strategy S2. Similarly, we observe that a more secure setting increases the protection rate and decreases the attack success rates. However, the attacker would leave numbers of overwritten FBC in the memory pool in this strategy making the protection rate higher.
@@
5.3. Protection rates with heap spray
Heap spraying boosts attackers’ chance of overwriting a sensitive field in a victim object and hence, increases the attack success rate and decreases the protection rate. (see detailed strategies and formal analysis in § D.4 and § D.5). We use the default settings of S2malloc and evaluate how both attack and defense rates change of strategies (S1 and S2) with different numbers of sprayed objects and object sizes. 2(e) and 2(f) show the protection and attack success rates of reusing the same pointer (strategy S1-spray). 2(g) and 2(h) show the protection and attack success rates using a fresh pointer (strategy S2-spray).
@@ 5.3. Protection rates with heap spray
5.4. Illustrate the protection rates
@@
Table 2. Protection and attack success rates of attack rounds in mRuby issue 4001 using the two strategies.
5.5. Defending against real-world CVEs
@@
We pick six UAF-write vulnerabilities (seven exploits), that are more severe and of the most interest to us, and only one UAF-Read only CVE (CVE-2022-22620) to showcase how S2malloc thwarts this attack. Results are summarized in Table 3. S2malloc is the only memory allocator that can thwart UAF-write attacks. Guarder can detect three exploits by recognizing double-free attempts. For UAF-read attacks, S2malloc uses RIO to stop the attacker from accurately reusing the memory chunk. DieHarder zeros out the memory after it is freed, which is effective only if the attacker is trying to over-read a freed block.
Vulnerability Attack Guarder DieHarder SlimGuard S2malloc
CVE-2015-6831 Info leak Detect (end of exec.) No defense No defense Detect via FBC change
CVE-2015-6834 Info leak Detect (end of exec.) No defense No defense Detect via FBC change
CVE-2015-6835 Info leak Detect (end of exec.) No defense No defense Detect via FBC change
CVE-2015-6835 Info leak + sleep No defense No defense No defense Detect via FBC change
CVE-2020-24346 Info leak No defense No defense No defense Detect via FBC change
Python-91153 Info leak No defense No defense No defense Detect via FBC change
mruby-4001 Info leak No defense No defense No defense Detect via FBC change
CVE-2022-22620 Info leak No defense Defense (zero-out) No defense Non-deterministic leak (RIO)
Table 3. Summary of how different memory allocators defend against eight exploitation techniques on seven different vulnerabilities.
6. Performance Evaluation
@@
Experiment setup.Experiment setupExperiment setup. Experiments are performed on both x86 and AARCH servers for macro benchmarks. The performance of benchmarks is measured only on the x86 server. The x86 server is configured with 64-bit 160-core 2.40GHz Intel Xeon E7-8870 CPUs (x86 architecture) with 1TB system memory. We set up the AARCH server on Amazon Web Service (AWS), using the im4gn.4xlarge machine with 16 vCPU cores and 64 GB memory. On both machines, benchmarks are measured in the Docker environment with Debian 11, kernel version 5.15.0. We measure the overheads using the GNU time binary (Linux Foundation, 2023b) and setting the LD_PRELOAD environment variable to substitute the system default allocator.
@@
Run-time Overhead (std. *$*10^{-3}$) Memory Overhead (std. *$*10^{-3}$)
@@
S2malloc 12% (4.1) 2.8% (1.0) 16% (2.3) 1.8% (0.9) 37% (0.2) 27% (1.9) 38% (0.0) 28% (0.7)
SlimGuard 17% (7.6) 4.4% (1.5) 7.7% (6.9) 2.6% (1.3) 57% (3.8) 23% (2.2) 57% (2.1) 24% (0.2)
DieHarder 31% (2.1) 2.1% (1.3) - 2.5% (0.7) 59% (0.0) 21% (1.2) - 21% (0.8)
Guarder 3.5% (5.5) 2.4% (11) - - 56% (0.1) 58% (0.7) - -
Table 4. Normalized run time and memory overheads for state-of-the-art entropy-based secure allocators on SPEC and PARSEC benchmarks. We report geometric averages and standard deviations of the run-time overhead over five runs.
6.1. Macro benchmarks
@@
PARSEC.PARSECPARSEC. We first evaluate the performance of S2malloc using the PARSEC (Bienia et al., 2008) benchmark. We exclude three network tests (netdedup, netferret, and netstreamcluster) and one test (x264) that fails to compile in the baseline scenario, and only report the result of the rest 12 benchmarks. Additionally, we exclude “raytrace” from execution for the AARCH sever as it cannot compile. We refer to each PARSEC test using the first three letters of its name.
@@
SPEC CPU2017.SPEC CPU2017SPEC CPU2017. We use SPEC CPU2017 (Standard Performance Evaluation Corporation, [n. d.]) version 1.1.9. We report the results of 12 C/C++ only tests in both “Integer” and “Floating Point” test suites with the default OpenMP settings of four parallel threads. All reported SPEC overheads are “reportable” following the SPEC documentation (Standard Performance Evaluation Corporation, [n. d.]).
@@
Results.ResultsResults. We measure the performance of S2malloc and three other entropy-based allocators (SlimGuard, Dieharder and Guarder) on both benchmarks with the x86 machine. On the AARCH machine, we exclude Guarder from analysis, noticing that Guarder relies on AES-NI (Intel Corporation, [n. d.]), an Intel CPU extension, not supported on AARCH machines. We report the average and standard deviations of overheads in Table 4. Averages and standard deviations of each test in both settings are shown in Appendix B. Missing columns in the figures indicate the corresponding execution runs into error. A complete list of erroneous executions and explanations is listed in Appendix A.
@@
Table 5. Normalized run-time and memory overheads of running mimalloc-benchmark
@@
Table 5. Normalized run-time and memory overheads of running mimalloc-benchmark
Table 6. Normalized memory and run-time Overhead changes compared with the default Settings.
6.2. Micro benchmarks
Figure 4. Execution time of glibc-simple
Figure 4. Execution time of glibc-simple
Figure 5. Run-time and memory overheads of running PARSEC with multi-threads
To further understand the overheads introduced by S2malloc, we investigate its performance using mimalloc-bench (daanx, [n. d.]), composed of real-world and calibrated programs that allocate and free heap memory frequently. The results are shown in Table 6. Individual results are reported in Appendix C. Running three tests with SlimGuard never returns (marked as ”-”) in Appendix C. They are excluded from computing the SlimGuard average overheads.
@@ 6.2. Micro benchmarks
We take one test, glibc-simple from mimalloc-bench, from the glibc micro-benchmark suite (gli, [n. d.]) to further investigate the delays incurred in the two most common heap object management functions – malloc() and free(). The test times the execution of allocating and freeing a large number of blocks of a given size. We modify the benchmark and monitor the execution time of calling malloc() and free() separately. To investigate the time consumption of different sizes, we vary the block size S to be 16B, 128B, and 1KB, and change the number of allocated blocks N correspondingly, so that the total allocation size (N * S) is always 1000 MB. Results are presented in Figure 5, and is the average of 100 runs.
@@
6.3. Performance on real world programs
To evaluate the performance of S2malloc in real-world environments, we run two servers: Nginx (1.18.0), and Lighttpd (1.4.71), and two databases: Redis (7.2.1), and SQLite (3.25.2) on the x86 machine. We use ApacheBench (ab) (The Apache Software Foundation, 2023) 2.3 to test the throughput and delays using the Nginx default root page, of 613 bytes, as the requested page with 500 concurrent requests. On Redis, we use the same settings as its performance is measured in mimalloc-bench (daanx, [n. d.]). We use sqlite-bench (Hajime Tazaki, [n. d.]) to measure the performance of SQLite. We report the results of performing random read and write operations in Table 7 and Table 8. We observe that applying S2malloc on these programs results in minimal throughput influence (even better throughput on Nginx and Redis). Running S2malloc delays the request response time on Nginx but not on Lighttpd. Applying all secure memory allocators increases memory consumption.

### v2

@@ 4.4 Hardening heap canaries
4.5 Free block canaries (FBC)
@@
Initially in S2malloc, we create the memory pool using the mmap system call with the ANON flag and all allocated memories are set to zero in the Linux environment [22]. We take this advantage and use the zeros as the initial FBC with the following benefits:
@@ 4.6 Summary and comparison
Being an entropy-based allocator, S2malloc is inherently closer to this line of work [29, 33, 24] with a nearly identical set of heap exploitation protection features except UAF protection. S2malloc provides a much stronger security assurance in the presence of UAF vulnerabilities. In particular, S2malloc addresses the two entropy-loss cases (discussed in § 3.2 and § 3.3) with RIO (§ 4.2) and RBL (§ 4.3), respectively, and hence, providing much higher effectiveness on UAF mitigation. In addition, S2malloc is designed to actively monitor the integrity of the heap and watch for UAF attempts, including heap spraying practices that aim to prepare the heap data and layout for UAF exploits. S2malloc achieves this through a synergy of regular heap canaries (§ 4.4) and FBC (§ 4.5).
On the other hand, Table 1 also shows a sheer contrast between entropy-based allocators and UAF-mitigating allocators. Notably, although providing a theoretically complete mitigation guarantee toward UAF, UAF-mitigating allocators [3, 42] significantly impair program efficiency and are hard to be deployed in time-sensitive use cases. In contrast, as will be presented in § 6, S2malloc incurs a significantly lower overhead that is typical for entropy-based allocators, making S2malloc practical and deployable on production systems if the end-user can tolerate a marginal chance of protection failure (less than 10% in the default setting of S2malloc, discussed in § 5).
5 Security Evaluation
@@
Figure 3: Parameterized security evaluation (x-axis: logarithmic block size/byte, y-axis: protection/attack success rate/%).
@@ 5.1 Formal analysis
For each of the scenarios above, we discuss the rates of whether the attacker adopts the heap spray technique or not separately. We introduce a total of four scenarios. We refer readers to Appendix 0.E for a detailed explanation of them.
5.2 Parameterized protection rates
@@
3(a) shows how the protection rate changes with a different set of tunable parameters while 3(b) shows how the attack success rate changes using strategy S1. We observe that adopting a more secure setting increases the protection rate and decreases the attack success rates for both small and large blocks. However, as the block size increases, both rates decrease as it is less likely to overwrite either the target field or FBC. Both the attacker and the defender are likely not to succeed, and the rates are thus not summed up to one (in fact, not even close to one).
3(c) shows how the protection rate changes with the same set of tunable parameters while 3(d) shows how the attack success rate changes using strategy S2. Similarly, we observe that a more secure setting increases the protection rate and decreases the attack success rates. However, the attacker would leave numbers of overwritten FBC in the memory pool in this strategy making the protection rate higher.
@@
5.3 Protection rates with heap spray
Heap spraying boosts attackers’ chance of overwriting a sensitive field in a victim object and hence, increases the attack success rate and decreases the protection rate. (see detailed strategies and formal analysis in § 0.E.4 and § 0.E.5). We use the default settings of S2malloc and evaluate how both attack and defense rates change of strategies (S1 and S2) with different numbers of sprayed objects and object sizes. 3(e) and 3(f) show the protection and attack success rates of reusing the same pointer (strategy S1-spray). 3(g) and 3(h) show the protection and attack success rates using a fresh pointer (strategy S2-spray).
@@ 5.3 Protection rates with heap spray
5.4 Illustrate the protection rates
@@
Table 2: Protection and attack success rates of attack rounds in mRuby issue 4001 using the two strategies.
5.5 Defending against real-world CVEs
@@
1typedef union _zvalue_value {
2    long lval;
3    double dval;
4    struct { char *val; int len; } str;
5    HashTable *ht;
6    zend_object_value obj;
7} zvalue_value;
8
9typedef struct _zval_struct {
10    zvalue_value value;
11    zend_uint refcount;
12    zend_uchar type;
13    zend_uchar is_ref;
14} zval;
Figure 4: Type definition of zval.
Vulnerability Attack pattern [33] [29] [24] S2malloc [19] [4]
CVE-2015-6831 DP $\rightarrow$ Write
CVE-2015-6834 DP $\rightarrow$ Write
CVE-2015-6835 DP $\rightarrow$ Write
CVE-2015-6835 DP $\rightarrow$ Write $\rightarrow$ sleep
CVE-2020-24346 DP $\rightarrow$ Write
Python-91153 DP $\rightarrow$ Write $\blacksquare$
mruby-4001 DP $\rightarrow$ Write $\blacksquare$ $\blacksquare$
CVE-2022-22620 DP $\rightarrow$ Read
[…]
$\bullet$
Table 3: Sumsmary of how different memory allocators defend against eight exploitation techniques on seven vulnerabilities. Vanilla BIBOP allocator and Scudo [25] are vulnerable to all attacks and behave similarly to Guarder [33] (DP: dangling pointer, ○: no defense, ◐: detect at the end of execution, ◑: defense via zero-out, ●: detect via FBC change,
[…]
$\bullet$
: non-deterministic leak (RIO), $\blacksquare$: thwart the exploitation ability).
Out of the seven vulnerabilities, six are UAF-write bugs and one (CVE-2022-22620) is a UAF-read only bug. We also found seven exploits against these vulnerabilities (two exploits for CVE-2015-6835 with different attack patterns). All CVEs except Python-91153 and mruby-4001 can cause arbitrary code execution (ACE) if properly exploited. However, a powerful attack (e.g., ACE by overriding a function pointer) can succeed when the target object is precisely allocated to a freed memory chunk that is still referred to by a dangling pointer [^1] An attacker may attack blindly, e.g., overriding a code pointer through the dangling pointer regardless of whether it points to a target object or not. This will have three consequences: 1) corrupting FBC, which may cause the attack to be detected upon future mallocs, 2) overwriting a wrong field due to RIO which may cause the program to enter a weird state (e.g., crash), 3) a successful attack. If the program can be recovered from a weird state automatically (e.g., crash resilience), the attacker can retry the same attack and eventually case 1 or case 3 will occur. However, without the probabilistic detection on UAF attempts by S2malloc, only case 3 will occur.. S2malloc can mitigate attacks by reducing the chance that a target object is referred to by a dangling pointer. Evaluation results of the eight exploits are in Table 3.
[…]
@@ 5    HashTable *ht;
Entropy-based allocators.Entropy-based allocatorsEntropy-based allocators. S2malloc can thwart all UAF-write attacks evaluated, Guarder can detect three exploits by recognizing double-free attempts, but DieHarder and SlimGuard fail to thwart these exploits. For the UAF-read attack, S2malloc uses RIO to stop the attacker from reusing the memory chunk with accurate object alignment, causing the data read by the dangling pointer to be not sensible. DieHarder zeros out the memory after it is freed, which is effective if the attacker tries to over-read a freed block.
[…]
Context-based allocators.Context-based allocatorsContext-based allocators. While we expect context-based allocators to demonstrate strong and stable protection, some of the exploits, unfortunately, hit on certain blind spots in Cling and TypeAfterType by accident.
In the case of Cling, if both the dangling pointer and the target object (i.e., the object an attacker hope to corrupt) are allocated through the same multi-layer function call sequence, they are considered to fall under the same allocation context, causing the target object to be possibly accessed by the dangling pointer. We will illustrate this limitation through the CVE-2015-6835 case study presented later. In fact, all examined CVEs, except mruby-4001, hit this limitation of Cling. Cling mitigates mruby-4001 by limiting the attacker to target objects of type mrb_io, which prevents the attacker from creating a powerful attack primitive.
TypeAfterType can unpack malloc wrappers with an arbitrary number of layers until it finds a sizeof(T) in the function argument, and an ID i is given to each allocation site of T. The tuple (i,T) makes the allocation context, and all memory allocations through this call sequence will be allocated from a memory pool dedicated to this context. However, if the dangling pointer and the target object share the same context in an exploit, UAF can still occur. We will illustrate this limitation through the CVE-2015-6835 case study. TypeAfterType mitigates Python-91153 by limiting the target object to be a reallocated string, and mruby-4001 by limiting the target object to be an mrb_io.
[…]
Case study: CVE-2015-6835.Case study: CVE-2015-6835Case study: CVE-2015-6835. This CVE is a UAF bug in the PHP session deserializer, which reconstructs a session from a serialized string. (Figure 7). An attacker can exploit this vulnerability to control a dangling pointer to a freed zval object. This is possible as the return value (a zval pointer) phpvarunserialize is freed in its caller without noticing that the same pointer might also be stored in a global variable SESSIONVARS.
The zval type, unfortunately, is a reference-counting wrapper over nearly all other objects involved in the PHP engine (see definition of zval in Figure 4). Therefore, an attacker might corrupt any zval object that may be reallocated to the freed slot. They can simply uses the echo(..) function to dump a newly allocated zval in the freed memory.
1) Protection from entropy-based allocators. S2malloc checks FBC on every malloc(). In this exploit, when the attacker tries to use the dangling pointer in zendechohandler, its refcount field is increased, causing the FBC to be modified. This enables S2malloc to detect the UAF attempt when the corrupted slot or a nearby slot is about to be reallocated. If the refcount change does not corrupt FBC (simulated by disabling the FBC check) and this corrupted block is reallocated, S2malloc can still stop the exploit as RIO causes misalignment between the dangling pointer and new object, causing the type field to have value UNKNOWN and prevents echo printing.
Guarder and DieHarder try to mitigate this attack by random allocation: hoping the new object will not be referred to by a dangling pointer. However, our experiment shows that Guarder fails if the attacker re-runs the attack multiple times or spray the heap with victim objects. SlimGuard fails to provide protections as it always allocates the most recently freed objects to the program. It does not implement the claimed random allocation feature and does not have any other security features that could detect UAF. DieHarder zeros out the memory chunk that stops information leakage of the freed zval, but it cannot prevent an attacker to corrupt the newly allocated zval over the freed chunk.
[…]
For Cling, this malloc wrapper implies that the allocation of many zval objects will be sharing the same context (measured by the two innermost return addresses on the call stack). This leaves the dangling pointer plenty of candidate objects to refer to after several rounds of deserialization in PHP. TypeAfterType can inline malloc wrappers but the inlining stops at phpvarunserialize because it sees the sizeof(zval) argument in emalloc and hence, will allocate all zval objects originating from this malloc wrapper from the same pool. Unfortunately, the dangling pointer is also allocated this way, enabling UAF among the dangling pointer to other zval objects as well.
[…]
Summary.SummarySummary. The combination of random allocation and delayed free-list provided by previous entropy-based allocators focus on one-time attacks only, Hence, repeating the same attack is a simple yet effective solution to undermine their protection. Context-based allocators, on the other hand, might fail to detect UAF among objects allocated of the same context. While these results highlight the effectiveness of S2malloc, we note that information leakage through corrupted pointers might diminish this guarantee, which is discussed in § 7.
6 Performance Evaluation
@@
Experiment setup.Experiment setupExperiment setup. Experiments are performed on both x86 and AARCH servers for macro benchmarks. The performance of benchmarks is measured only on the x86 server. The x86 server is configured with 64-bit 160-core 2.40GHz Intel Xeon E7-8870 CPUs (x86 architecture) with 1TB system memory. We set up the AARCH server on Amazon Web Service (AWS), using the im4gn.4xlarge machine with 16 vCPU cores and 64 GB memory. On both machines, benchmarks are measured in the Docker environment with Debian 11, kernel version 5.15.0. We measure the overheads using the GNU time binary [23] and setting the LD_PRELOAD environment variable to substitute the system default allocator.
@@
Run-time Overhead Memory Overhead
@@
S2malloc 12% 2.8% 16% 1.8% 37% 27% 38% 28%
SlimGuard 17% 4.4% 7.7% 2.6% 57% 23% 57% 24%
DieHarder 31% 2.1% - 2.5% 59% 21% - 21%
Guarder 3.5% 2.4% - - 56% 58% - -
Table 4: Normalized overheads on SPEC and PARSEC benchmarks.
6.1 Macro benchmarks
@@
PARSEC.PARSECPARSEC. We first evaluate the performance of S2malloc using the PARSEC [6] benchmark. We exclude three network tests (netdedup, netferret, and netstreamcluster) and one test (x264) that fails to compile in the baseline scenario, and only report the result of the rest 12 benchmarks. Additionally, we exclude “raytrace” from execution for the AARCH sever as it cannot compile. We refer to each PARSEC test using the first three letters of its name.
@@
SPEC CPU2017.SPEC CPU2017SPEC CPU2017. We use SPEC CPU2017 [36] version 1.1.9. We report the results of 12 C/C++ only tests in both “Integer” and “Floating Point” test suites with the default OpenMP settings of four parallel threads. All reported SPEC overheads are “reportable” following the SPEC documentation [36].
@@
Results.ResultsResults. We measure the performance of S2malloc and three other entropy-based allocators (SlimGuard, Dieharder and Guarder) on both benchmarks with the x86 machine. On the AARCH machine, we exclude Guarder from analysis, noticing that Guarder relies on AES-NI [18], an Intel CPU extension, not supported on AARCH machines. We report the average and standard deviations of overheads in Table 4. Averages and standard deviations of each test in both settings are shown in Appendix 0.B. Missing columns in the figures indicate the corresponding execution runs into error. A complete list of erroneous executions and explanations is listed in Appendix 0.A.
@@
Table 5: Normalized run-time and memory overheads of running mimalloc-benchmark
@@
Table 6: Normalized memory and run-time Overhead changes compared with the default Settings.
6.2 Micro benchmarks
Figure 5: Execution time of glibc-simple
Figure 6: Run-time and memory overheads of running PARSEC with multi-threads
To further understand the overheads introduced by S2malloc, we investigate its performance using mimalloc-bench [9], composed of real-world and calibrated programs that allocate and free heap memory frequently. The results are shown in Table 6. Individual results are reported in Appendix 0.C. Running three tests with SlimGuard never returns (marked as ”-”) in Appendix 0.C. They are excluded from computing the SlimGuard average overheads.
@@ 6.2 Micro benchmarks
We take one test, glibc-simple from mimalloc-bench, from the glibc micro-benchmark suite [1] to further investigate the delays incurred in the two most common heap object management functions – malloc() and free(). The test times the execution of allocating and freeing a large number of blocks of a given size. We modify the benchmark and monitor the execution time of calling malloc() and free() separately. To investigate the time consumption of different sizes, we vary the block size S to be 16B, 128B, and 1KB, and change the number of allocated blocks N correspondingly, so that the total allocation size (N * S) is always 1000 MB. Results are presented in Figure 6, and is the average of 100 runs.
@@
6.3 Performance on real world programs
To evaluate the performance of S2malloc in real-world environments, we run two servers: Nginx (1.18.0), and Lighttpd (1.4.71), and two databases: Redis (7.2.1), and SQLite (3.25.2) on the x86 machine. We use ApacheBench (ab) [37] 2.3 to test the throughput and delays using the Nginx default root page, of 613 bytes, as the requested page with 500 concurrent requests. On Redis, we use the same settings as its performance is measured in mimalloc-bench [9]. We use sqlite-bench [2] to measure the performance of SQLite. We report the results of performing random read and write operations in Table 7 and Table 8. We observe that applying S2malloc on these programs results in minimal throughput influence (even better throughput on Nginx and Redis). Running S2malloc delays the request response time on Nginx but not on Lighttpd. Applying all secure memory allocators increases memory consumption.